Please help in showing me how to achieve the correct answer for the following:
The U.S. Senate consists of 100 members, 2 from each state. A committee of 9 senators is formed. What is the probability that it contains at least one senator from your state? (Round your answer to two decimal places.)

Answer by stanbon(75887)

Please help in showing me how to achieve the correct answer for the following:
The U.S. Senate consists of 100 members, 2 from each state. A committee of 9 senators is formed. What is the probability that it contains at least one senator from your state? (Round your answer to two decimal places.)
---
P(at least one) = 1 - P(none from my state)
-----
My state has 2 of the 100.
Ans: 1 - 98C2/100C2 = 1 - 0.9602 = 0.0398
----------------------
Cheers,
Stan H.
----------------------